Between which two consecutive integers is root 239 located?
Answer by fcabanski(1391)

To solve this problem without a calculator guess a number that, when squared, = 239. Try 13: 13 * 13 = 169. That's too small. 14 * 14 = 196. Too small. 15 * 15 = 225. And 16*16 = 256. So the root of 239 (when a number isn't specified, root means square root) is between 15 and 16.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)